OpenStudy (anonymous):

How do you represent the (decimal) integer 50 in, oh, "hexadecimal," otherwise known as base-16? Recall that decimal is simply base-10, and binary is simply base-2. Infer from those base systems how to represent this one! *

OpenStudy (anonymous):

Decimal is converted into hexadecimal by dividing it by 16 decimal/16 = hexadecimal 50/16 = 32

OpenStudy (e.mccormick):

Just remember the extra didgits: Decimal: 0123456789 Hex: 0123456789ABCDEF So you have 16 hex spots and the last is F. That is when you need a second column, when you pass F. Like how we have 9, 10, 11 in decimal there is F, 10, 11 in hex. Is it just they do not mean the same values. Here, I'll char some values to show what I mean. \(\begin{array}{ccc} \text{Dec} & = & \text{Hex}\\ 1 & = & 1\\ 5 & = & 5\\ 10 & = & A\\ 11 & = & B\\ 12 & = & C\\ 13 & = & D\\ 14 & = & E\\ 15 & = & F\\ 16 & = & 10\\ 17 & = & 11\\ 18 & = & 12\\ 26 & = & 1A\\ 27 & = & 1B\\ 128 & = & 80\\ 192 & = & C0\\ 219 & = & DB\\ 229 & = & E5\\ \end{array}\)

OpenStudy (mandre):

The easiest way to convert a number from decimal to another form is to divide by the base you want. For example to convert 19 to binary 19 / 2 = 9 remainder 1 9 / 2 = 4 remainder 1 4 / 2 = 2 remainder 0 2 / 2 = 1 remainder 0 1 /2 = 0 remainder 1 You are done once you get to 0 and your binary number will be the remainder from bottom to top: 10011 To do Hex 50 / 16 = 3 remainder 2 3 / 16 = 0 remainder 3 So 50 converted to Hexadecimal is 32.

OpenStudy (e.mccormick):

And to show how what Mandre did works with what I did, look at this one: 26 / 16 = 1 remainder 10 1 / 16 = 0 remainder 1 Well, in HEX, 10 is A. So this is really: 26 / 16 = 1 remainder A 1 / 16 = 0 remainder 1 26 dec = 1A hex. That means it still works and all you really need to remember is how to find what A, B, C, D, E, and F are.